### v1

@@ Abstract
The solution of sequences of shifted linear systems is a classic problem in numerical linear algebra, and a variety of efficient methods have been proposed over the years. Nevertheless, there still exist challenging scenarios witnessing a lack of performing solvers. For instance, state-of-the-art procedures struggle to handle nonsymmetric problems where the shifts are complex numbers that do not come as conjugate pairs. We design a novel projection strategy based on the rational Krylov subspace equipped with a minimal residual condition. We also devise a novel pole selection procedure, tailored to our problem, providing poles for the rational Krylov basis construction that yield faster convergence than those computed by available general-purpose schemes. A panel of diverse numerical experiments shows that our novel approach performs better than state-of-the-art techniques, especially on the very challenging problems mentioned above.
@@
Shifted linear systems, Sylvester matrix equation, rational Krylov method, minimal residual condition.
@@ 1 Introduction
with $A\in\mathbb{R}^{n\times n}$, $I_{n}$ the identity matrix of order $n$, $s_{i}$ a possibly complex number, and $b^{(i)}\in\mathbb{R}^{n}$ is a classic problem in numerical linear algebra. For the sake of simplicity in the presentation, we assume the right-hand side in (1) to be independent of the parameter index $i$, namely $b\equiv b^{(i)}$ for all $i=1,\ldots,\ell$. However, many of the results presented in this talk can be extended to a more general framework; see section 4.
Equations (1) can be met in numerous application settings, such as control theory and model order reduction techniques [4], structural dynamics [9], lattice quantum-chromodynamics [13], Tikhonov regularization [12], the solution of time-dependent PDEs [19], and the inner step in sophisticated eigensolvers [25], to name a few.
The ubiquity of (1) has driven the interest of our community into developing many different strategies for the solution of this problem over the years; see, e.g., the survey [35] and the references therein. Most of these techniques build upon the construction of the polynomial Krylov subspace
(2) $\mathbf{K}_{m}(A,r_{0})=\text{span}\{r_{0},Ar_{0},\ldots,A^{m-1}r_{0}\},$
where $r_{0}=b-Ax_{0}$ is the initial residual vector for a given initial guess $x_{0}\in\mathbb{R}^{n}$. In these schemes, solutions of the form
(3) $x_{m}^{(i)}=x_{0}+V_{m}y_{m}^{(i)},$
are constructed where the columns of $V_{m}=[v_{1},\ldots,v_{m}]\in\mathbb{R}^{n\times m}$ form an orthonormal basis of (2) whereas $y_{m}^{(i)}\in\mathbb{C}^{m}$. The backbone of polynomial Krylov subspace methods for shifted linear systems is the shift-invariant property of (2), namely the following Arnoldi relation
(4) $(A+s_{i}I_{n})V_{m}=V_{m}(H_{m}+s_{i}I_{m})+h_{m+1,m}v_{m+1}e_{m}^{T}=V_{m+1}\left(\underline{H}_{m}+s_{i}\begin{bmatrix}I_{m}\\ 0\\ \end{bmatrix}\right),$
holds true for any shift $s_{i}$ in (1). In (4), $H_{m}\in\mathbb{R}^{m\times m}$ and $h_{m+1,m}\in\mathbb{R}$ are quantities related to the orthonormalization step of the Arnoldi process and $\underline{H}_{m}=[H_{m};h_{m+1,m}e_{m}^{T}]\in\mathbb{R}^{(m+1)\times m}$. The relation (4) explains the use of the shift-independent space (2) employed in the construction of the solutions (3). Indeed, (4) allows us to construct a single basis $V_{m}$ and deal with shifted quantities only at the projected level.
@@
$(H_{m}+s_{i}I)y_{m}^{(i)}=\beta e_{1},\quad\beta=\|r_{0}\|;$
see, e.g., [27].
@@
$y_{m}^{(i)}=\operatorname*{arg\,min}_{y\in\mathbb{C}^{m}}\|r_{0}-(A+s_{i}I)V_{m}y\|=\operatorname*{arg\,min}_{y\in\mathbb{C}^{m}}\left\|\beta e_{1}-\left(\underline{H}_{m}+s_{i}\begin{bmatrix}I_{m}\\ 0\\ \end{bmatrix}\right)y\right\|,$
see, e.g., [27].
In many cases, polynomial Krylov subspace methods for (1) require the construction of a large subspace to meet the prescribed level of accuracy. This leads to severe drawbacks in terms of both computational cost and storage demand, especially for those methods where a full orthogonalization step and the allocation of the whole $V_{m}$ are necessary. Different strategies have been developed to overcome these issues. Since preconditioning is often not an option as designing good preconditioners for all the shifted linear systems in the sequence (1) is rather difficult, the restarting paradigm has become one of the most popular approaches in this setting; see, e.g., [30, 11, 10]. However, it is well-known that restarting needs collinearity of the residual vectors $r_{m}^{(i)}$ to be successful and this feature may be difficult to achieve in minimal residual methods unless rather strict conditions on $A$ and the shifts $s_{i}$ are considered; see, e.g., [10].
Even though they have been designed for solving more general problems like general parametrized linear systems, recycling techniques can also be employed for the solution of shifted linear systems. The main underlying idea here is to solve a given problem, e.g., by a Krylov subspace method, and reuse the most significant part of this approximation space in the solution of another problem. See, e.g., [33, 36, 14, 1] for more details on recycling techniques.
In principle, one could combine restart and recycling. However, this integration is not always possible due to the need for residual collinearity to restart; see, e.g., [35, Section 7.1.1] and [36, Theorem 1].
A completely different approach consists in employing approximation spaces different from (2) in the construction of the solutions (3). This novel point of view has been proposed for the first time in [32] where tools coming from the matrix equation literature have been adapted to the solution of (1). Indeed, it is easy to show that the sequence of shifted linear systems in (1) with $b\equiv b^{(i)}$, for all $i$, can be reformulated in terms of a single Sylvester matrix equation of the form
@@
In this setting, an approximate solution to (5) of the form $X_{m}=V_{m}Y_{m}$ is sought. In [32], the columns of $V_{m}\in\mathbb{R}^{n\times 2m}$ represent an orthonormal basis for the extended Krylov subspace
@@
$T_{m}Y_{m}+Y_{m}S=\beta e_{1}\mathbf{1}_{\ell}^{T},\quad T_{m}=V_{m}^{T}AV_{m},\;\beta=\|b\|.$
In [32] it has been shown that the use of the extended Krylov subspace (6) in place of its polynomial counterpart (2) as approximation space leads to a much faster convergence in terms of number of iterations so that restarting is not really necessary. On the other hand, the computation of $V_{m}$ is now significantly more expensive than before due to solving linear systems with $A$ at each iteration.
Our novel method takes inspiration from what has been proposed in [32] but differs in both the selection of the approximation space and the computation of the matrix $Y_{m}$. In our scheme, we assume that we can afford solving a handful of shifted linear systems by, e.g., a sparse direct solver, but that we do not have the computational resources for solving all the $\ell$ linear systems in (1). This can be the case when, e.g., we need to solve (1) on a laptop where we may be able to solve $(A+sI)x=b$ but the lack of a proper parallel computing environment makes the sequential solution of all the systems in (1) extremely time-consuming.
@@
(7) $\mathbf{K}_{m}(A,b,\bm{\xi}_{m-1})=\text{span}\left\{b,(A+\xi_{2}I)^{-1}b,\ldots,\prod_{i=2}^{m}(A+\xi_{i}I)^{-1}b\right\},$
@@
in place of (6) as approximation space. In (7), $\bm{\xi}_{m-1}=(\xi_{2},\ldots,\xi_{m})^{T}\in\mathbb{C}^{m-1}$ is a set of poles that can be either given a priori or computed on the fly as the space expands. Various strategies for the computation of these poles are available in the literature; see, e.g., [6, 15]. Selecting good poles is crucial to obtain a performing rational Krylov subspace method for the problem at hand. In the following, we will illustrate an ad-hoc routine taking into account the peculiar structure of our problem (5). Numerical experiments show that our approach often performs better than off-the-shelf pole-selection strategies.
@@
Moreover, we compute $Y_{m}\in\mathbb{C}^{m\times\ell}$ by imposing a minimal residual condition instead of a Galerkin one as done in [32]. In contrast to what happens when (2) is adopted as approximation space, this appealing minimization property does not come with the drawback of forcing collinearity of the residuals, which may not even exist. Indeed, in our extensive numerical experience, the rational Krylov subspace one needs to construct is always of very moderate dimension, so that no restart has to be performed.
@@
Here is a synopsis of the paper. In section 2, we recall some results depicting scenarios where we can indeed expect the solution $X$ to (5) to be numerically low rank. This property is key for the success of projection-based solvers for matrix equations. Section 3 presents the main contribution of this paper, namely the derivation of a novel minimal residual rational Krylov subspace method for (1). As already mentioned, the performance of this method is related to the quality of the poles selected for the construction of the space (7). In section 3.1, an effective pole-selection strategy tailored to the problem (1) is presented. The construction of the rational Krylov subspace (7) requires the solution of shifted linear systems with $A$. At first, we derive our new numerical scheme by assuming that we are able to solve these linear systems by a direct method. However, this may not always be the case. Therefore, in section 3.2 we discuss a fully iterative variant of our minimal residual rational Krylov scheme, where the matrix $A$ is required only in the computation of matrix-vector products. The numerical experiments presented in section 5 show how our novel solver outperforms state-of-the-art techniques for shifted linear systems on very demanding problems like those where, e.g., the shifts $s_{i}$’s are complex and do not come in complex conjugate pairs. The paper ends with some conclusions in section 6.
@@ 2 When to expect a low-rank solution
Intuition suggests that whenever the shifts $s_{i}$’s are all “close” to each other, then the solutions $x^{(i)}$ will be so as well, leading to a matrix $X=[x^{(1)},\ldots,x^{(\ell)}]$ of low rank. To the best of our knowledge, this has never been formalized in the literature.
A different tool to understand when our strategy may be successful is the convergence theory behind rational Krylov methods for Sylvester equations; see, e.g., [2]. However, many of these results rely on the field of values of the coefficient matrices overlooking the diagonal pattern of $S$ in our case.
@@
$\Lambda\widetilde{X}+\widetilde{X}S=\widetilde{b}\mathbf{1}_{\ell}^{T},$
@@
Since both $\Lambda$ and $S$ are diagonal, $\widetilde{X}$ is a Cauchy matrix of the form $(\widetilde{X})_{i,j}=\widetilde{b}_{i}/(\lambda_{i}+s_{j})$. Several results on the decay of the singular values of Cauchy matrices can be found in the literature. In particular, the rate of this decay is often bounded from above by quantities that depend on the distance between the $s_{i}$’s and the $\lambda_{j}$’s; see, e.g., [3, Section 4.1][37, Theorem 3.2], and [38, Eq. (2.34)]. None of this contribution takes into account the possible clustering of the shifts $s_{i}$’s.
In the following proposition, we derive a novel result which relates the low-rank approximability of the Cauchy matrix $\widetilde{X}$ to the relative distance between the shifts $s_{i}$’s and a “centroid” $\mathbf{s}$, and the distance between the shifts and the eigenvalues of $A$.
@@
Assume that there exists a centroid $\mathbf{s}$ such that $\max_{i=1,\ldots,\ell}|s_{i}-\mathbf{s}|\leq\epsilon$ and that $\min_{j=1,\ldots,n}|\lambda_{j}+\mathbf{s}|/\epsilon\gg 1$. Then
(8) $\|\widetilde{X}-(\Lambda+\mathbf{s}I)^{-1}\widetilde{b}\mathbf{1}_{\ell}^{T}\|_{F}\leq c(\epsilon,\mathbf{s},\Lambda)\|\widetilde{X}\|_{F},$
where $c(\epsilon,\mathbf{s},\Lambda)=\epsilon/\min_{j=1,\ldots,n}|\lambda_{j}+\mathbf{s}|\ll 1$.
@@
The $(i,j)$-th entry of $\widetilde{X}$ is of the form $(\widetilde{X})_{i,j}=\widetilde{b}_{i}/(\lambda_{i}+s_{j})$. We thus have
@@
The result follows by taking the square root of the summation over the indices $i$ and $j$ of the square of the terms above.
The assumptions of Proposition 2.1 portray a scenario where the shifts $s_{i}$’s are much closer to the centroid $\mathbf{s}$ than to the mirrored eigenvalues $-\lambda_{j}$’s, underlying the importance of having well-conditioned problems. Indeed, having close shifts could not be sufficient to obtain (almost) linearly dependent solutions if the related shifted linear systems are badly conditioned.
Proposition 2.1 can be easily generalized to the case where we can identify a certain number $m$ of different clusters of shifts, each of them equipped with its centroid $\mathbf{s}_{k}$, $k=1,\ldots,m$. In particular, by applying the result of Proposition 2.1 column-wise, cluster by cluster, we can show that $\widetilde{X}$ can be well approximated by a rank-$m$ matrix.
@@
We generate the matrix $A\in\mathbb{R}^{n\times n}$, $n=200$, as $A=Q\Lambda Q^{-1}$, where $Q\in\mathbb{R}^{n\times n}$ is a matrix with random entries drawn from the Gaussian distribution, whereas $\Lambda=\text{diag}(-100,\ldots,-1,1,\ldots,100)$. We consider $\ell=300$ complex shifts of the form $s_{j}=0.5+i\omega_{j}10^{-k}$, $k>0$, with $\omega_{j}$ a random scalar drawn once again from the Gaussian distribution. We can think of the exponent $k$ as an index of the level of clustering of the shifts: larger $k$ yields more clustered $s_{j}$’s. We select $b\in\mathbb{R}^{n}$ as a random vector with random entries drawn from the Gaussian distribution.
@@
Table 1: Example 2.3. $A=Q\Lambda Q^{-1}\in\mathbb{R}^{n\times n}$, $n=200$, $Q\in\mathbb{R}^{n\times n}$ random matrix with entries drawn from the Gaussian distribution, $\Lambda=\text{diag}(-100,\ldots,-1,1,\ldots,100)$. $b\in\mathbb{R}^{n}$ random vector. We consider two different sets of $\ell=300$ shifts. Left: $s_{j}=0.5+i\omega_{j}10^{-k}$, $j=1,\ldots,300$. Right: $s_{j}=1+10^{-10}+i\omega_{j}10^{-k}$, $j=1,\ldots,300$. $k$ denotes the level of clustering of the constructed shifts. $X\in\mathbb{C}^{n\times\ell}$ denotes the exact solution computed column-wise by “backslash”.
@@
Note that regardless of $k$, the minimum distance between the shifts and the eigenvalues of $A$ is 0.5 and that the shifts lie within the convex hull of the spectrum of $A$. Therefore, state-of-the-art results solely relying on the distance between the $s_{j}$’s and the eigenvalues of $A$ cannot predict how the low-rank approximability of $X$ depends on the clustering of the shifts, as illustrated by the results in Table 1 (left). On the other hand, this aspect is accounted for by Proposition 2.1. Indeed, from the way we constructed the shifts, we have $\mathbf{s}=0.5$ and $\max_{i=1,\ldots,\ell}|s_{i}-\mathbf{s}|=\mathcal{O}(10^{-k})$ so that $c(\epsilon,\mathbf{s},\Lambda)=\mathcal{O}(10^{1-k})$.

### v2

@@ Abstract
The solution of sequences of shifted linear systems is a classic problem in numerical linear algebra, and a variety of efficient methods have been proposed over the years. Nevertheless, there exist challenging scenarios for which there is still a lack of performant solvers. For instance, state-of-the-art procedures based on restarting may struggle to handle nonsymmetric problems where the shifts are possibly complex numbers that do not come as conjugate pairs. We design a novel projection strategy based on the rational Krylov subspace equipped with a minimal residual condition. We also devise a novel pole selection procedure, tailored to our problem, providing poles for the rational Krylov basis construction that yield faster convergence than those computed by available general-purpose schemes. A panel of diverse numerical experiments shows that our novel approach performs better than state-of-the-art techniques, especially on the very challenging problems mentioned above.
@@
Shifted linear systems, Sylvester matrix equations, rational Krylov method, minimal residual condition.
@@ 1 Introduction
with $A\in{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbb{C}}^{n\times n}$, $I_{n}$ the identity matrix of order $n$, $s_{i}$ a possibly complex number, and $b^{(i)}\in\mathbb{R}^{n}$ is a classic problem in numerical linear algebra. For the sake of simplicity in the presentation, we assume the right-hand side in (1) to be independent of the parameter index $i$, namely $b\equiv b^{(i)}$ for all $i=1,\ldots,\ell$. However, many of the results presented in this paper can be extended to a more general framework; see section 4.
Equations (1) is encountered in numerous application settings, such as control theory and model order reduction techniques [MOR2015], structural dynamics [FERIANI2000], lattice quantum-chromodynamics [Frommer1], Tikhonov regularization [Frommer2], the solution of time-dependent PDEs [TimeDepPDES], and the inner step in sophisticated eigensolvers [FEAST], to name a few.
The ubiquity of (1) has driven the interest of our community into developing many different strategies for the solution of this problem over the years; see, e.g., the survey [Survey] and the references therein. Most of these techniques build upon the construction of the polynomial Krylov subspace
(2) $\mathbf{K}_{m}(A,b)=\text{span}\{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}b},A{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}b},\ldots,A^{m-1}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}b}\}.$
In these schemes, solutions of the form
(3) $x_{m}^{(i)}=V_{m}y_{m}^{(i)},$
are constructed where the columns of $V_{m}=[v_{1},\ldots,v_{m}]\in\mathbb{C}^{n\times m}$ form an orthonormal basis of (2) whereas $y_{m}^{(i)}\in\mathbb{C}^{m}$. The backbone of polynomial Krylov subspace methods for shifted linear systems is the shift-invariance property of (2), namely the following Arnoldi relation
[…]
holds true for any shift $s_{i}$ in (1). The underlined matrix $\underline{X}$ in this work denotes a matrix formed by concatenating a square matrix $X$ with an additional last row or a block row for block methods. In particular, in (4), $H_{m}\in{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbb{C}}^{m\times m}$ and $h_{m+1,m}\in{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbb{C}}$ are quantities related to the orthonormalization step of the Arnoldi process and $\underline{H}_{m}=[H_{m};h_{m+1,m}e_{m}^{T}]\in{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbb{C}}^{(m+1)\times m}$. The relation (4) explains the use of the shift-independent space (2) employed in the construction of the solutions (3). Indeed, (4) allows us to construct a single basis $V_{m}$ and deal with shifted quantities only at the projected level.
@@
$(H_{m}+s_{i}I)y_{m}^{(i)}=\beta e_{1},\quad\beta=\|{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}b}\|;$
see, e.g., [Saad2003].
@@
$y_{m}^{(i)}=\operatorname*{arg\,min}_{y\in\mathbb{C}^{m}}\|{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}b}-(A+s_{i}I)V_{m}y\|=\operatorname*{arg\,min}_{y\in\mathbb{C}^{m}}\left\|\beta e_{1}-\left(\underline{H}_{m}+s_{i}\begin{bmatrix}I_{m}\\ 0\\ \end{bmatrix}\right)y\right\|,$
see, e.g., [Saad2003].
In many cases, polynomial Krylov subspace methods for (1) require the construction of a large subspace to meet the prescribed level of accuracy. This leads to severe drawbacks in terms of both computational cost and storage demand, especially for those methods where a full orthogonalization step and the allocation of the whole $V_{m}$ are necessary. Different strategies have been developed to overcome these issues. Even though some contributions aiming to construct effective preconditioning operators for (1) can be found in the literature (see, e.g., [Bau2015, Saibaba2013, Bakhos2017]), preconditioning is seldom an option as designing good preconditioners for all the shifted linear systems in the sequence (1) is rather difficult. Therefore, the restarting paradigm is often the preferred option in this setting; see, e.g., [FOM, RestartedGMRES, BicgStab]. However, it is well-known that restarting needs collinearity of the residual vectors $r_{m}^{(i)}$ to be successful and this feature may be difficult to achieve in minimal residual methods unless rather strict conditions on $A$ and the shifts $s_{i}$ are considered; see, e.g., [BicgStab].
Even though they have been designed for solving more general problems like general parametrized linear systems, recycling techniques can also be employed for the solution of shifted linear systems. The main underlying idea here is to solve a given problem, e.g., by a Krylov subspace method, and reuse the most significant part of this approximation space in the solution of another problem. See, e.g., [Rec1, rec2, Gaul2014, AldGHR21] for more details on recycling techniques.
In principle, one could combine restart and recycling. However, this integration is not always possible due to the need for residual collinearity to restart; see, e.g., [Survey, Section 7.1.1] and [rec2, Theorem 1].
A completely different approach consists of employing approximation spaces different from (2) in the construction of the solutions (3). This novel point of view has been proposed for the first time in [Simoncini2010] where tools coming from the matrix equation literature have been adapted to the solution of (1). Indeed, it is easy to show that the sequence of shifted linear systems in (1) with $b\equiv b^{(i)}$, for all $i$, can be reformulated in terms of a single Sylvester matrix equation of the form
@@
In this setting, an approximate solution to (5) of the form $X_{m}=V_{m}Y_{m}$ is sought. In [Simoncini2010], the columns of $V_{m}\in{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbb{C}}^{n\times 2m}$ represent an orthonormal basis for the extended Krylov subspace
@@
$T_{m}Y_{m}+Y_{m}S=\beta e_{1}\mathbf{1}_{\ell}^{T},\quad T_{m}=V_{m}^{*}AV_{m},\;\beta=\|b\|.$
In [Simoncini2010] it has been demonstrated that the use of the extended Krylov subspace (6) in place of its polynomial counterpart (2) as approximation space leads to a much faster convergence in terms of number of iterations so that restarting is not really necessary. However, the computation of $V_{m}$ is now significantly more expensive than before due to solving linear systems with $A$ at each iteration.
Our novel method takes inspiration from what has been proposed in [Simoncini2010] but differs in both the selection of the approximation space and the computation of the matrix $Y_{m}$. In our scheme, we assume that we can afford solving a handful of shifted linear systems by, e.g., a sparse direct solver, but that we do not have enough resources for solving all the $\ell$ linear systems in (1). This can be the case when, e.g., we need to solve (1) on a laptop where we may be able to solve $(A+sI)x=b$ but the lack of a proper parallel computing environment makes the sequential solution of all the systems in (1) extremely time-consuming, especially for large $\ell$. This issue may be overcome by working with HPC infrastructures. However, the parallel solution of $\ell$ large-scale linear systems is still an extremely energy-draining operation so that being able to solve much fewer linear systems can remarkably lower the carbon footprint of the overall solution procedure.
@@
(7) $\mathbf{K}_{m+1}(A,b,\bm{\xi}_{m})=\text{span}\left\{b,(A+\xi_{1}I)^{-1}b,\ldots,\prod_{i=1}^{m}(A+\xi_{i}I)^{-1}b\right\},$
@@
in place of (6) as approximation space. In (7), $\bm{\xi}_{m}=(\xi_{1},\ldots,\xi_{m})^{T}\in\mathbb{C}^{m}$ is a set of poles that can be either given a priori or computed on the fly as the space expands. Various strategies for the computation of these poles are available in the literature; see, e.g., [Poles1, Gue13b]. Selecting good poles is crucial to obtain a performant rational Krylov subspace method for the problem at hand. In section 3.1, we illustrate an ad-hoc greedy routine taking into account the peculiar structure of our problem (5). Numerical experiments show that our approach often performs better than off-the-shelf pole-selection strategies.
The rational Krylov relation that holds at each iteration of the rational Arnoldi procedure [blockrationalKrylov] to construct $V_{m+1}$, the orthonormal basis for $\mathbf{K}_{m+1}(A,b,\bm{\xi}_{m})$, is
(8) $AV_{m+1}\underline{K}_{m}=V_{m+1}\underline{H}_{m},$
@@
where $\underline{K}_{m}$ and $\underline{H}_{m}$ are $(m+1)\times m$ unreduced Hessenberg matrices that result from the orthonormalization procedure. In this work we propose to seek a solution under the form $X_{m}=V_{m+1}\underline{K}_{m}Y_{m}$, where we compute $Y_{m}\in\mathbb{C}^{m\times\ell}$ by imposing a minimal residual condition instead of a Galerkin one as done in [Simoncini2010]. In contrast to what happens when (2) is adopted as approximation space, this appealing minimization property does not come with the drawback of forcing collinearity of the residuals, which may not even exist. Indeed, in our extensive numerical experience, the rational Krylov subspace one needs to construct is always of very moderate dimension, so that no restart has to be performed.
In [GuZhouLin2007] the authors proposed to employ flexible GMRES for (1), preconditioned with operators of the form $\mathcal{P}_{j}=A+\alpha_{j}I$ for a small number of predefined scalars $\alpha_{j}$. Even though this can be viewed as a first, timid step towards our rational Krylov subspace method, the storage of the preconditioned and unpreconditioned bases in flexible strategies makes the need of restarting even more urgent; this carries over all the difficulties related to restarting we mentioned above.
We would like to mention that a peculiar rational Krylov subspace is adopted in [illposed] for the solution of ill-posed problems. In particular, the authors propose to construct what they call a mixed space that we can formally view as (7) where at every other iteration an infinite pole is employed, i.e., $\xi_{2k}=\infty$, $k\geq 1$, whereas the poles with odd index are possibly good regularization parameters chosen from a given set. Even though the ultimate goal in [illposed] is rather different from ours, we believe that the rational Krylov approach presented here may be seen as a generalization of the mixed approach from [illposed] when it comes to Tikhonov regularization. The interesting extension of our approach to the solution of ill-posed problems will be studied elsewhere in the near future.
@@
Here is a synopsis of the paper. In section 2, we recall some results depicting scenarios in which we can indeed expect the solution $X$ to (5) to be numerically low rank. This property is key for the success of projection-based solvers for matrix equations. Section 3 contains the main contribution of this paper, namely the derivation of a novel minimal residual rational Krylov subspace method for (1). As already mentioned, the performance of this method is related to the quality of the poles selected for the construction of the space (7). In section 3.1, an effective pole-selection strategy tailored to the problem (1) is presented. The construction of the rational Krylov subspace (7) requires the solution of shifted linear systems with $A$. At first, we derive our new numerical scheme by assuming that we are able to solve these linear systems by a direct method. However, this may not always be the case. Therefore, in section 3.2 we discuss a fully iterative variant of our minimal residual rational Krylov scheme, where the matrix $A$ is required only in the computation of matrix-vector products. In section 4 we generalize our algorithm to problems (1) whose right-hand sides $b^{(i)}$’s are not fixed. This is attained by using block rational Krylov subspaces. The numerical experiments presented in section 5 show how our novel solver outperforms state-of-the-art techniques for shifted linear systems on very demanding problems like those where, e.g., the shifts $s_{i}$’s are complex and do not come in complex conjugate pairs. The paper ends with some conclusions in section 6.
@@ 2 When to expect a low-rank solution
Intuition suggests that whenever the shifts $s_{i}$’s are all “close” to each other, then the solutions $x^{(i)}$ will be as well, leading to a matrix $X=[x^{(1)},\ldots,x^{(\ell)}]$ of low rank. To the best of our knowledge, this property has been exploited only in [Grasedyck04] where Grasedyck shows the decay of the singular values of the solution $X$ to a general Sylvester equation AXXB=CAX-XB=C assuming the eigenvalues of $B$ to be clustered. On the other hand, the techniques emloyed to get these results require $\text{Re}(\lambda-\gamma)<0$ for all eigenvalues $\lambda$ of $A$ and $\gamma$ of $B$. This may be a too restrictive hypothesis in case of (5).
A different tool to understand when our strategy may be successful is the convergence theory behind rational Krylov methods for Sylvester equations; see, e.g., [ConvRKSylv]. However, many of these results rely on the field of values of the coefficient matrices overlooking the diagonal pattern of $S$ in our case.
@@
(9) $\Lambda\widetilde{X}+\widetilde{X}S=\widetilde{b}\mathbf{1}_{\ell}^{T},$
@@
Since both $\Lambda$ and $S$ are diagonal, $\widetilde{X}$ is a Cauchy matrix of the form $(\widetilde{X})_{i,j}=\widetilde{b}_{i}/(\lambda_{i}+s_{j})$. Several results on the decay of the singular values of Cauchy matrices can be found in the literature. In particular, the rate of this decay is often bounded from above by quantities that depend on the distance between the $s_{i}$’s and the mirrored $\lambda_{j}$’s; see, e.g., [BecT17, Section 4.2][Truhar1, Theorem 3.2], and [Truhar2, Eq. (2.34)]. None of this contribution takes into account the possible clustering of the shifts $s_{i}$’s.
In the following proposition, we derive a novel result which relates the low-rank approximability of the solution $X$ to the relative distance between the shifts $s_{i}$’s and a “centroid” $\mathbf{s}$, and the distance between the shifts and the eigenvalues of $A$.
@@
Let $A=Q\Lambda Q^{-1}$, $\Lambda=\text{diag}(\lambda_{1},\ldots,\lambda_{n})$, and assume that there exists a centroid $\mathbf{s}$ such that $\max_{i=1,\ldots,\ell}|s_{i}-\mathbf{s}|\leq\epsilon$ and that $\min_{j=1,\ldots,n}|\lambda_{j}+\mathbf{s}|/\epsilon\gg 1$. Then
(10) $\|X-(A+\mathbf{s}I)^{-1}b\mathbf{1}_{\ell}^{T}\|_{F}\leq c(\epsilon,\mathbf{s},\Lambda)\kappa_{F}(Q)\|X\|_{F},$
where $c(\epsilon,\mathbf{s},\Lambda)=\epsilon/\min_{j=1,\ldots,n}|\lambda_{j}+\mathbf{s}|\ll 1$ and $\kappa_{F}(Q)=\|Q\|_{F}\|Q^{-1}\|_{F}$.
@@
Let’s first consider the transformed equation (9). The $(i,j)$-th entry of $\widetilde{X}$ is of the form $(\widetilde{X})_{i,j}=\widetilde{b}_{i}/(\lambda_{i}+s_{j})$. We thus have
@@
By taking the square root of the summation over the indices $i$ and $j$ of the square of the terms above, it is then easy to show that
(11) $\|\widetilde{X}-(\Lambda+\mathbf{s}I)^{-1}\widetilde{b}\mathbf{1}_{\ell}^{T}\|_{F}\leq c(\epsilon,\mathbf{s},\Lambda)\|\widetilde{X}\|_{F}.$
We thus have
$\|X-(A+\mathbf{s}I)^{-1}b\mathbf{1}_{\ell}^{T}\|_{F}\leq\|Q\|_{F}\|\widetilde{X}-(\Lambda+\mathbf{s}I)^{-1}\widetilde{b}\mathbf{1}_{\ell}^{T}\|_{F}\leq c(\epsilon,\mathbf{s},\Lambda)\|Q\|_{F}\|\widetilde{X}\|_{F},$
and the result follows by recalling that $\widetilde{X}=Q^{-1}X$.
The assumptions of Proposition 2.1 portray a scenario where the shifts $s_{i}$’s are much closer to the centroid $\mathbf{s}$ than to the mirrored eigenvalues $\lambda_{j}$’s, underlying the importance of having well-conditioned problems. Indeed, having close shifts could not be sufficient to obtain (almost) linearly dependent solutions if the related shifted linear systems are badly conditioned.
Proposition 2.1 can be easily generalized to the case where we can identify a certain number $m$ of different clusters of shifts, each of them equipped with its centroid $\mathbf{s}_{k}$, $k=1,\ldots,m$. In particular, by applying the result of Proposition 2.1 column-wise, cluster by cluster, we can show that $X$ can be well approximated by a rank-$m$ matrix.
@@
We generate the matrix $A\in\mathbb{R}^{n\times n}$, $n=200$, as $A=Q\Lambda Q^{-1}$, where $Q\in\mathbb{R}^{n\times n}$ is a matrix with random entries drawn from the Gaussian distribution***rng(’default’), $\kappa_{F}=\mathcal{O}(10^{3})$, whereas $\Lambda=\text{diag}(-100,\ldots,-1,1,\ldots,100)$. We consider $\ell=300$ complex shifts of the form $s_{j}=0.5+i\omega_{j}10^{-k}$, $k>0$, with $\omega_{j}$ a random scalar drawn once again from the Gaussian distribution. We can think of the exponent $k$ as an index of the level of clustering of the shifts: larger $k$ yields more clustered $s_{j}$’s. We select $b\in\mathbb{R}^{n}$ as a random vector with random entries drawn from the Gaussian distribution and unit norm.
@@
Table 1: Example 2.3. $A=Q\Lambda Q^{-1}\in\mathbb{R}^{n\times n}$, $n=200$, $Q\in\mathbb{R}^{n\times n}$ random matrix with entries drawn from the Gaussian distribution, $\kappa_{F}=\mathcal{O}(10^{3})$, $\Lambda=\text{diag}(-100,\ldots,-1,1,\ldots,100)$. $b\in\mathbb{R}^{n}$ random vector. We consider two different sets of $\ell=300$ shifts. Left: $s_{j}=0.5+i\omega_{j}10^{-k}$, $j=1,\ldots,300$. Right: $s_{j}=1+10^{-10}+i\omega_{j}10^{-k}$, $j=1,\ldots,300$. $k$ denotes the level of clustering of the constructed shifts. $X\in\mathbb{C}^{n\times\ell}$ denotes the exact solution computed column-wise by “backslash”.
@@
Note that regardless of $k$, the minimum distance between the shifts and the eigenvalues of $A$ is 0.5 and that the shifts lie within the field of values of $A$. Moreover, $\text{Re}(\lambda+s_{j})$ does not have a constant sign for all eigenvalues $\lambda$ of $A$ and shifts $s_{j}$, but it can be either positive or negative. Therefore, to the best of our knowledge, state-of-the-art results cannot predict how the low-rank approximability of $X$ depends on the clustering of the shifts, as illustrated by the results in Table 1 (left). This aspect is accounted for by Proposition 2.1. Indeed, from the way we constructed the shifts, we have $\mathbf{s}=0.5$ and $\max_{i=1,\ldots,\ell}|s_{i}-\mathbf{s}|=\mathcal{O}(10^{-k})$ so that $c(\epsilon,\mathbf{s},\Lambda)=\mathcal{O}(10^{1-k})$.